i forgot the last digit of a 7 digit telephone number.if i randomly dial the final 3 digits after correctly diagonising the first four,then what is the chance of dialing the correct answer

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Aparna
 
It is given that last three digits are randomly dialled. then each of the digit can be selected out of 10 digits in 10 ways.
Hence required probability =(1/10)³ = 1/1000
 
Regards
Arun (askIITians forum expert)